how many liters of 80% acid solution must be mixed with 15% acid solution to get 455 L of a 70% acid solution?
Answer by lwsshak3(11628)

how many liters of 80% acid solution must be mixed with 15% acid solution to get 455 L of a 70% acid solution?
***
let x=amt of 80% solution to mix
455-x=amt of 15% solution to mix
80%x+15%(455-x)=70%*455
.80x+68.25-.15x=318.5
0.65x=250.25
x=385
455-x=70
amt of 80% solution to mix=385 liters
amt of 15% solution to mix=70 liters